# Privately Evaluating Untrusted Black-Box Functions

### v1

@@ 3.2 Sens-o-Matic: Removing the Monotonicity Promise
Fix a universe $\mathcal{U}$ and a range $\mathcal{Y}\subseteq\mathbb{R}$. For each $\ell\in\mathbb{Z}$, the $-\ell$ monotonization of a function $f:\mathcal{U}^{*}\to\mathcal{Y}$ is the function ${\text{\sf M}}_{\ell}[{f}]:\mathcal{U}^{*}\to\mathcal{Y}$ defined by777 We use the convention that if $\mathcal{Y}$ is unbounded below, then $\inf(\mathcal{Y})=-\infty$.
@@
4:Set $\lambda\leftarrow 2\cdot\lambda_{\text{ShI}}(\varepsilon/2,\delta,\beta/2,|\mathcal{Y}|)$, where $\lambda_{\text{ShI}}$ is given by Lemma 3.2. $\triangleright$ $\lambda$ is set so that ShI run with the parameter settings below uses depth parameter $\lambda/2$.
@@
The two failure events we consider are (1) when the noise variable $Z$ in 5 has large absolute value, $|Z|>\frac{2}{\varepsilon}\ln\frac{2}{\beta}$, and (2) when ShI fails. Each of these events happens with probability at most $\beta/2$, by 2.2 and the setting of parameters given to ShI. By the union bound over these two events, the overall failure probability is at most $\beta.$
@@
$\mathcal{W}^{f}(x)\geq\min_{x^{\prime}\in{\mathcal{N}}^{\downarrow}_{\lambda/2}(x)}{\text{\sf M}}_{\ell}[{f}](x^{\prime})\geq\min_{x^{\prime\prime}\in{\mathcal{N}}^{\downarrow}_{\lambda}(x)}f(x^{\prime\prime})=f(x)-DS_{\lambda}^{f}(x).$
@@ 4 Privacy Wrappers with Claimed Sensitivity Bound
Theorem 4.1 (Subset-Extension privacy wrapper).
There exists a constant $a>0$ such that for every universe $\mathcal{U}$, privacy parameters $\varepsilon>0,\delta\in(0,1)$, and Lipschitz constant $c>0$, there exists an $(\varepsilon,\delta)$-privacy wrapper $\mathcal{W}$ over $\mathcal{U}$ with noise distribution $\text{Lap}{\left({\frac{a\cdot c}{\varepsilon}}\right)}$ for all $c$-Lipschitz functions $f:\mathcal{U}^{*}\to\mathbb{R}$ and all $x\in\mathcal{U}^{*}$. Moreover, $\mathcal{W}$ is $O{\left({\frac{1}{\varepsilon}\log\frac{1}{\delta}+1}\right)}$-down local for all $x\in\mathcal{U}^{*}$.
@@ 5 Locality Lower Bound
Our next theorem states that the query complexity of any privacy wrapper that achieves the same accuracy guarantee as that of Theorem 3.1 must depend on the cardinality of the range of the function.
@@ 6 Query Complexity Lower Bound
    If $\frac{1}{\varepsilon}\log\min(\frac{r}{\alpha\beta},\frac{1}{\delta})\leq r\leq n^{.49}$ then $q=n^{\Omega\left(\tfrac{1}{\varepsilon}\log\min\left(\tfrac{r}{\alpha\beta},\tfrac{1}{\delta}\right)\right)}$.
@@
    If $r\leq\min(\frac{1}{\varepsilon}\log\min(\frac{r}{\alpha\beta},\frac{1}{\delta}),n^{.49})$ then $q=n^{\Omega(r)}$.
@@
A lower bound on query complexity directly implies a lower bound on locality, since a $\lambda$-down local algorithm makes at most $\binom{|x|}{\lambda}$ distinct queries. However, the locality lower bound implied by Theorem 6.1 is weaker than Theorem 5.1 in two ways. First, for the setting of large $r$ and $\delta=0$, the locality lower bound implied by Theorem 6.1 is at most $\frac{1}{\varepsilon}\log\frac{n}{\alpha\beta}$, while the locality lower bound given by Theorem 5.1 is $\frac{1}{\varepsilon}\log\frac{r}{\alpha\beta}$. Second, by Theorem 5.3, in the automated sensitivity detection setting the locality must have at least $\log^{*}$ dependence on the size of the range. While we also deduce a query complexity lower bound in this setting (Remark 6.2), the locality lower bound it implies is weaker than the bound given by Theorem 5.3, since the query complexity lower bound may not have any dependence on the size of the range.
Tightness of our bounds
Below we discuss implications of Theorem 6.1 to the tightness of our upper bounds in both the automated sensitivity detection setting and in the claimed sensitivity bound setting.
@@
Since all Lipschitz functions $f$ have $DS^{f}_{\alpha}(x)\leq\alpha$ for all datasets $x$, a privacy wrapper that is $(DS^{f}_{\alpha}(x),\beta)$-accurate for all functions $f$ and datasets $x$, is also $(\alpha,\beta)$-accurate for all Lipschitz functions $f$ and datasets $x$. It follows that Theorem 6.1 also holds for privacy wrappers that are $(DS^{f}_{\alpha}(x),\beta)$-accurate for all functions $f:\mathcal{U}^{*}\to\{0,1,\dots,r\}$. Recall that Theorem 3.1 gives a privacy wrapper for the automated sensitivity detection setting that has query complexity $n^{\lambda(\varepsilon,\delta,\beta,r)}$. In the setting where $\delta=0$, Item 1 of Theorem 6.1 implies that the query complexity of this privacy wrapper cannot be improved. In the setting of $\delta>0$, the query complexity of our privacy wrapper differs from the lower bound in Item 1 of Theorem 6.1 by a small factor of $2^{O{\left({\log^{*}r}\right)}}$ in the exponent.
@@
Plugging in parameters to Lemma 6.4, we prove each item of Theorem 6.1. To prove Item 1 we set $\rho=r$, then, since $\rho\leq n^{.49}$ we see that $\kappa=\frac{1}{\varepsilon}\log\min(\frac{\rho}{\alpha\beta},\frac{\varepsilon}{\delta})$, and thus $\frac{n}{\rho\kappa}=n^{\Theta(1)}$, which completes the proof of Item 1. To prove Item 2 we set $\rho=r$, then, since $r\leq\min(\frac{1}{\varepsilon}\log\min(\frac{r}{\alpha\beta},\frac{1}{\delta}),n^{.49})$, we see that $\kappa=\rho$ and $\frac{n}{\rho\kappa}=n^{\Theta(1)}$, which completes the proof of Item 2. Last, to prove Item 3 we set $\alpha=\varepsilon n$ and $\rho=3\varepsilon n$, observe that this implies $\kappa=\frac{n}{2\rho}=\Omega(\frac{1}{\varepsilon})$, and $\frac{n}{\rho\kappa}=2$, completing the proof of Item 3. ∎
@@ Acknowledgments
We are grateful to Jonathan Ullman for helpful conversations and discussion of our results, and notably their extension to , as well as to
@@ Appendix B Utility Analysis of Our Version of Kohli-Laskowski’s TAHOE
For every universe $\mathcal{U}$, privacy parameters $\varepsilon>0,\delta\in(0,1)$, and Lipschitz constant $c>0$, there exists an $(\varepsilon,\delta)$-privacy wrapper $\mathcal{W}$ over $\mathcal{U}$ with noise distribution $\text{Lap}{\left({\frac{36c}{\varepsilon^{2}}\ln\frac{1}{\delta}}\right)}$ for all $c$-Lipschitz functions $f:\mathcal{U}^{*}\to\mathbb{R}$ and all $x\in\mathcal{U}^{*}$. Moreover, $\mathcal{W}$ is $O{\left({\frac{1}{\varepsilon}\log\frac{1}{\delta}}\right)}$-down local and has query complexity $|x|^{O{\left({\frac{1}{\varepsilon}\log\frac{1}{\delta}}\right)}}$ for all $x\in\mathcal{U}^{*}$.
@@
4:set $\varepsilon_{0}\leftarrow\frac{\varepsilon}{3}$ and $\delta_{0}\leftarrow\delta/3e^{\varepsilon_{0}}$ and $\tau\leftarrow\frac{1}{\varepsilon_{0}}\ln\frac{1}{\delta_{0}}$
5:release $\ell\leftarrow|x|-5\tau-r_{1}$ where $r_{1}\sim\text{TruncLap}{\left({\frac{1}{\varepsilon_{0}},\tau}\right)}$triangleright Definition 2.7
6:$h\leftarrow|x|-\tau-r_{2}$ where $r_{2}\sim\text{TruncLap}{\left({\frac{1}{\varepsilon_{0}},\tau}\right)}$triangleright $h$ is not released
@@
8:else return $f(u)+Z$ where $u=\arg\max\{|v|:v\in\Sigma^{f}_{{\ell},{h}}(x)\}$, and $Z\sim\text{Lap}{\left({\frac{12\tau}{\varepsilon_{0}}}\right)}$triangleright If more than one such $u$ exists then pick one arbitrarily
@@
Let $\mathcal{W}$ denote Algorithm 3, and fix neighbors $u,v\in\mathcal{U}^{*}$. We first argue that $\mathcal{W}$ is $(\varepsilon,\delta)$-DP. For each $\ell\in\mathbb{Z}$, let ${\mathcal{A}}_{\ell}$ denote the steps of $\mathcal{W}$ after releasing $\ell$. Our analysis is split into two major steps. First, we argue that for fixed $\ell$, algorithm ${\mathcal{A}}_{\ell}$ is DP, then, we argue that $\mathcal{W}$ releases $\ell$ privately. DP composition (2.3) then suffices to complete the proof.
Observe that for $\ell\in[\min(|u|,|v|)-6\tau,\min(|u|,|v|)-4\tau]$, if neither ${\mathcal{A}}_{\ell}(u)$ nor ${\mathcal{A}}_{\ell}(v)$ output $\perp$, then there exists $u^{\prime}\subseteq u$ and $v^{\prime}\subseteq v$ such that $u^{\prime}$ and $v^{\prime}$ are $\ell$-stable with respect to $f$, and $|u^{\prime}\cap v^{\prime}|\geq\ell$. Thus, $|f(u^{\prime})-f(v^{\prime})|\leq|u\setminus(u^{\prime}\cap v^{\prime})|+|v\setminus(u^{\prime}\cap v^{\prime})|\leq 12\tau$. Hence, by the Laplace mechanism (2.1), for all measurable sets $E\subseteq\mathbb{R}\cup\{\perp\}$, and all $\ell\in[\min(|u|,|v|)-6\tau,\min(|u|,|v|)-4\tau]$,
$\Pr[{\mathcal{A}}_{\ell}^{f}(u)\in E|{\mathcal{A}}_{\ell}^{f}(u)\neq\perp]\leq e^{\varepsilon_{0}}\Pr[{\mathcal{A}}_{\ell}^{f}(v)\in E|{\mathcal{A}}_{\ell}^{f}(v)\neq\perp].$
Next, we argue that $\Pr[{\mathcal{A}}_{\ell}^{f}(u)\neq\perp]\leq e^{\varepsilon_{0}}\Pr[{\mathcal{A}}_{\ell}^{f}(v)\neq\perp]+\delta_{0}$. Let $m_{\ell}(x)=\max\{|v|:v\in\Sigma^{f}_{{\ell},{h}}(x)\}$. By 4.4, we have $|m_{\ell}(u)-m_{\ell}(v)|\leq 1$. Since ${\mathcal{A}}_{\ell}(u)$ outputs $\perp$ if and only if $\Sigma^{f}_{{\ell},{h}}(u)$ is empty, we need only analyze the probability that $h>m_{\ell}(u)$. Since $m_{\ell}$ has sensitivity $1$, and $h$ is distributed according to $\text{TruncLap}(\frac{1}{\varepsilon_{0}},\tau)$, the Laplace mechanism guarantees that $\Pr[{\mathcal{A}}_{\ell}^{f}(u)\neq\perp]\leq e^{\varepsilon_{0}}\Pr[{\mathcal{A}}_{\ell}^{f}(v)\neq\perp]+\delta_{0}$. Putting it all together,
$\Pr[{\mathcal{A}}^{f}_{\ell}(u)\in E]$ $=\Pr[{\mathcal{A}}_{\ell}^{f}(u)\in E|{\mathcal{A}}_{\ell}^{f}(u)\neq\perp]\Pr[{\mathcal{A}}_{\ell}^{f}(u)\neq\perp]+\Pr[{\mathcal{A}}_{\ell}^{f}(u)\in E|{\mathcal{A}}_{\ell}^{f}(u)=\perp]\Pr[{\mathcal{A}}_{\ell}^{f}(u)=\perp]$
$\leq e^{\varepsilon_{0}}\Pr[{\mathcal{A}}_{\ell}^{f}(v)\in E|{\mathcal{A}}_{\ell}^{f}(v)\neq\perp]{\left({e^{\varepsilon_{0}}\Pr[{\mathcal{A}}_{\ell}^{f}(v)\neq\perp]+\delta_{0}}\right)}$
$+\Pr[{\mathcal{A}}_{\ell}^{f}(v)\in E|{\mathcal{A}}_{\ell}^{f}(v)=\perp]{\left({e^{\varepsilon_{0}}\Pr[{\mathcal{A}}_{\ell}^{f}(v)=\perp]+\delta_{0}}\right)}$
$\leq e^{2\varepsilon_{0}}\Pr[{\mathcal{A}}^{f}_{\ell}(v)\in E]+2e^{\varepsilon_{0}}\delta_{0}.$
Since $||u|-|v||=1$, the Laplace mechanism guarantees that the “release $\ell$” step of $\mathcal{W}$ satisfies $(\varepsilon_{0},\delta_{0})$-DP. Thus, composition (2.3) implies that $\mathcal{W}^{f}$ is $(3\varepsilon_{0},3e^{\varepsilon_{0}}\delta_{0})$-DP, which, by our choice of $\varepsilon_{0}$ and $\delta_{0}$, completes the proof of privacy.
@@
To see why the accuracy guarantee holds, observe that if $x$ is $\ell$-stable with respect to $f$, then $x=\arg\max\{|v|:v\in\Sigma^{f}_{{\ell},{h}}(x)\}$, and hence $\mathcal{W}^{f}(x)$ outputs $f(x)+\text{Lap}{\left({\frac{12\tau}{\varepsilon_{0}}}\right)}$. ∎

### v2

@@ 3.2 Sens-o-Matic: A Wrapper for General Functions
Fix a universe $\mathcal{U}$ and a range $\mathcal{Y}\subseteq\mathbb{R}$. For each $\ell\in\mathbb{Z}$, the $-\ell$ monotonization of a function $f:\mathcal{U}^{*}\to\mathcal{Y}$ is the function ${\text{\sf M}}_{\ell}[{f}]:\mathcal{U}^{*}\to\mathcal{Y}$ defined by666 We use the convention that if $\mathcal{Y}$ is unbounded below, then $\inf(\mathcal{Y})=-\infty$.
@@
4:Set $\lambda\leftarrow 2\cdot\lambda_{\text{ShI}}(\varepsilon/2,\delta,\beta/2,|\mathcal{Y}|)$, where $\lambda_{\text{ShI}}$ is given by Lemma 3.2.
$\triangleright$ $\lambda$ is set so that ShI run with the parameter settings below uses depth parameter $\lambda/2$.
@@
The two failure events we consider are (1) the noise variable $Z$ in 5 has large absolute value, $|Z|>\frac{2}{\varepsilon}\ln\frac{2}{\beta}$, and (2) ShI fails. Each of these events happens with probability at most $\beta/2$, by 2.2 and the setting of parameters given to ShI. By the union bound over these two events, the overall failure probability is at most $\beta.$
@@
$\mathcal{W}^{f}(x)$ $\geq\min_{x^{\prime}\in{\mathcal{N}}^{\downarrow}_{\lambda/2}(x)}{\text{\sf M}}_{\ell}[{f}](x^{\prime})$
$\geq\min_{x^{\prime\prime}\in{\mathcal{N}}^{\downarrow}_{\lambda}(x)}f(x^{\prime\prime})\qquad=\quad f(x)-DS_{\lambda}^{f}(x).$
@@ 4 Privacy Wrappers with Claimed Sensitivity Bound
Theorem 4.1Subset extension privacy wrapper. (Subset-Extension privacy wrapper).
There exists a constant $a>0$ such that for every universe $\mathcal{U}$, privacy parameters $\varepsilon>0,\delta\in(0,1)$, and Lipschitz constant $c>0$, there exists an $(\varepsilon,\delta)$-privacy wrapper $\mathcal{W}$ over $\mathcal{U}$ with noise distribution $\text{Lap}{\left({\frac{a\cdot c}{\varepsilon}}\right)}$ for all $c$-Lipschitz functions $f:\mathcal{U}^{*}\to\mathbb{R}$ and all $x\in\mathcal{U}^{*}$. Moreover, $\mathcal{W}$ is $O{\left({\frac{1}{\varepsilon}\log\frac{1}{\delta}}\right)}$-down local for all $x\in\mathcal{U}^{*}$.
@@ 5 Locality Lower Bound
Our next theorem states that the locality of any privacy wrapper that achieves the same accuracy guarantee as that of Theorem 3.1 must depend on the cardinality of the range of the function.
@@ 6 Query Complexity Lower Bound
    If $\frac{1}{\varepsilon}\log\min(\frac{r}{\alpha\beta},\frac{1}{\delta})\leq r\leq n^{0.49}$ then $q=n^{\Omega\left(\tfrac{1}{\varepsilon}\log\min\left(\tfrac{r}{\alpha\beta},\tfrac{1}{\delta}\right)\right)}$.
@@
    If $r\leq\min(\frac{1}{\varepsilon}\log\min(\frac{r}{\alpha\beta},\frac{1}{\delta}),n^{0.49})$ then $q=n^{\Omega(r)}$.
@@
A lower bound on query complexity directly implies a lower bound on locality, since a $\lambda$-down local algorithm makes at most $\binom{|x|}{\lambda}$ distinct queries. However, the locality lower bound implied by Theorem 6.1 is weaker than Theorems 5.1 and 5.3—specifically, the locality bounds implied by Theorem 6.1 do not capture the correct dependence on the range size $r$. When $\delta=0$, the locality lower bound given by Theorem 5.1 is $\frac{1}{\varepsilon}\log\frac{r}{\alpha\beta}$, whereas the bound implied by Theorem 6.1 is at most $\frac{1}{\varepsilon}\log\frac{n}{\alpha\beta}$. When $\delta>0$, in the automated sensitivity detection setting, Theorem 5.3 implies that the locality must have at least $\log^{*}$ dependence on the range; in contrast, the locality lower bound implied by Theorem 6.1 has no dependence on the range.
@@
Since all Lipschitz functions $f$ have $DS^{f}_{\alpha}(x)\leq\alpha$ for all datasets $x$, a privacy wrapper that is $(DS^{f}_{\alpha}(x),\beta)$-accurate for all functions $f$ and datasets $x$ is also $(\alpha,\beta)$-accurate for all Lipschitz functions $f$ and datasets $x$. It follows that Theorem 6.1 also holds for privacy wrappers that are $(DS^{f}_{\alpha}(x),\beta)$-accurate for all functions $f:\mathcal{U}^{*}\to\{0,1,\dots,r\}$. Recall that Theorem 3.1 gives a privacy wrapper for the automated sensitivity detection setting that has query complexity $n^{\lambda(\varepsilon,\delta,\beta,r)}$. In the setting where $\delta=0$, Item 1 of Theorem 6.1 implies that the query complexity of this privacy wrapper cannot be improved. In the setting of $\delta>0$, the query complexity of our privacy wrapper differs from the lower bound in Item 1 of Theorem 6.1 by a factor of $2^{O{\left({\log^{*}r}\right)}}$.
@@
Plugging in parameters to Lemma 6.4, we prove each item of Theorem 6.1. To prove Item 1 we set $\rho=r$, then, since $\rho\leq n^{0.49}$ we see that $\kappa=\frac{1}{\varepsilon}\log\min(\frac{\rho}{\alpha\beta},\frac{\varepsilon}{\delta})$, and thus $\frac{n}{\rho\kappa}=n^{\Theta(1)}$, which completes the proof of Item 1. To prove Item 2 we set $\rho=r$, then, since $r\leq\min(\frac{1}{\varepsilon}\log\min(\frac{r}{\alpha\beta},\frac{1}{\delta}),n^{0.49})$, we see that $\kappa=\rho$ and $\frac{n}{\rho\kappa}=n^{\Theta(1)}$, which completes the proof of Item 2. Last, to prove Item 3 we set $\alpha=\varepsilon n$ and $\rho=3\varepsilon n$, observe that this implies $\kappa=\frac{n}{2\rho}=\Omega(\frac{1}{\varepsilon})$, and $\frac{n}{\rho\kappa}=2$, completing the proof of Item 3. ∎
@@ Acknowledgments
We are grateful to Jonathan Ullman for helpful conversations and discussion of our results, and notably their application to parameter estimation in Erdős–Rényi graphs.
@@ Appendix B Utility Analysis of Our Version of Kohli-Laskowski’s TAHOE
Let $a>0$ be a sufficiently large constant. For every universe $\mathcal{U}$, privacy parameters $\varepsilon>0,\delta\in(0,1)$, and Lipschitz constant $c>0$, there exists an $(\varepsilon,\delta)$-privacy wrapper $\mathcal{W}$ over $\mathcal{U}$ with noise distribution $\text{Lap}{\left({\frac{a\cdot c}{\varepsilon^{2}}\ln\frac{1}{\delta}}\right)}$ for all $c$-Lipschitz functions $f:\mathcal{U}^{*}\to\mathbb{R}$ and all $x\in\mathcal{U}^{*}$. Moreover, $\mathcal{W}$ is $O{\left({\frac{1}{\varepsilon}\log\frac{1}{\delta}}\right)}$-down local and has query complexity $|x|^{O{\left({\frac{1}{\varepsilon}\log\frac{1}{\delta}}\right)}}$ for all $x\in\mathcal{U}^{*}$.
@@
4:set $\varepsilon_{0}\leftarrow\frac{\varepsilon}{4}$ and $\delta_{0}\leftarrow\delta/3$ and $\tau\leftarrow\lceil\frac{1}{\varepsilon_{0}}\ln\frac{1}{\delta_{0}}\rceil$
5:release $\ell\leftarrow|x|-11\tau-r_{1}$ where $r_{1}\sim\text{TruncLap}{\left({\frac{1}{\varepsilon_{0}},\tau}\right)}$triangleright Definition 2.7
6:$h\leftarrow|x|-2\tau-r_{2}$ where $r_{2}\sim\text{TruncLap}{\left({\frac{2}{\varepsilon_{0}},2\tau}\right)}$triangleright $h$ is not released
@@
8:else return $f(u)+Z$ where $u=\arg\max\{|v|:v\in\Sigma^{f}_{{\ell},{|x|-4\tau}}(x)\}$, and $Z\sim\text{Lap}{\left({\frac{10\tau}{\varepsilon_{0}}}\right)}$triangleright If more than one such $u$ exists then pick one arbitrarily
@@
We will use the notation $\approx_{\varepsilon,\delta}$ from Definition 4.5, and $m_{\ell}$ from Definition 4.4 throughout the proof. Let $\mathcal{W}$ denote Algorithm 3. To analyze $\mathcal{W}$, it will be convenient to break the mechanism down in steps as in the proof of Theorem 4.1.
  1. 1.
    Let $\mathcal{L}(x)$ denote the mechanism which releases $\ell\leftarrow|x|-11\tau-r_{1}$ where $r_{1}\sim\text{TruncLap}{\left({\frac{1}{\varepsilon_{0}},\tau}\right)}$, and let ^$\widehat{\mathcal{L}}(x)$ denote the set of possible outputs of $\mathcal{L}(x)$.
Additionally, for all fixed $\ell\in\mathbb{Z}$,
  1. 1.
    Let ${\mathcal{T}}_{\ell}(x)$ denote the following mechanism: set $h\leftarrow|x|-2\tau-r_{2}$ where $r_{2}\sim\text{TruncLap}{\left({\frac{2}{\varepsilon_{0}},2\tau}\right)}$; return $b\leftarrow\mathds{1}{\left[{\Sigma^{f}_{{\ell},{h}}(x)\neq\emptyset}\right]}$.
  2. 2.
    Let ${\mathcal{A}}_{\ell}(x)$ be the mechanism which returns $g_{\ell}(x)+Z$ where $g_{\ell}(x)=f(u)$ for $u=\arg\max\{|u|:u\in\Sigma^{f}_{{\ell},{|x|-4\tau}}(x)\cup\emptyset\}$, and $Z\sim\text{Lap}{\left({\frac{10\tau}{\varepsilon_{0}}}\right)}$.
  3. 3.
    Let ${\mathcal{P}}_{\ell}(x)$ be the mechanism which runs ${\mathcal{T}}_{\ell}(x)$ and outputs $\perp$ if ${\mathcal{T}}_{\ell}(x)=0$ and outputs ${\mathcal{A}}_{\ell}(x)$ otherwise.
By inspection of Algorithm 3, one can easily see that $\mathcal{W}(x)$ is equivalent to the following mechanism: Set $\ell\leftarrow\mathcal{L}(x)$ and output ${\mathcal{P}}_{\ell}(x)$.
Lemma B.2 (Privacy for fixed $\ell$).
Fix neighbors $x,y\in\mathcal{U}^{*}$ and ^$\ell\in\widehat{\mathcal{L}}(x)\cup\widehat{\mathcal{L}}(y)$. Then ${\mathcal{P}}_{\ell}(x)\approx_{3\varepsilon_{0},2\delta_{0}}{\mathcal{P}}_{\ell}(y)$.
Proof.
We prove the lemma via the following two claims. For each $\ell\in\mathbb{Z}$, define the set of “good” points
$G_{\ell}={\left\{{(x,y)\colon m_{\ell}(x)\geq|x|-4\tau\wedge m_{\ell}(y)\geq|y|-4\tau}\right\}}.$
Claim B.3.
In the setting of Lemma B.2, we have ${\mathcal{T}}_{\ell}(x)\approx_{\varepsilon_{0},\delta_{0}}{\mathcal{T}}_{\ell}(y)$.
Proof.
Observe that ${\mathcal{T}}_{\ell}$ is a postprocessing of the mechanism which on input $|x|$ releases $m_{\ell}(x)-|x|+2\tau+r_{2}$ where $r_{2}\sim\text{TruncLap}{\left({\frac{2}{\varepsilon_{0}},2\tau}\right)}$. By 4.4 the function $m_{\ell}(\cdot)-|\cdot|$ has sensitivity at most $2$, and thus by the Laplace mechanism, ${\mathcal{T}}_{\ell}(x)\approx_{\varepsilon_{0},\delta_{0}}{\mathcal{T}}_{\ell}(y)$. ∎
Claim B.4.
In the setting of Lemma B.2, if $(x,y)\not\in G_{\ell}$ then $\Pr{\left[{{\mathcal{T}}_{\ell}(x)=1}\right]},\Pr{\left[{{\mathcal{T}}_{\ell}(y)=1}\right]}\leq\delta$.
Proof.
Suppose $x\subset y$. Then by 4.4 we have $m_{\ell}(y)\geq m_{\ell}(x)\geq m_{\ell}(y)-1$. If $m_{\ell}(x)<|x|-4\tau$ then $m_{\ell}(y)\leq m_{\ell}(x)+1<|y|-4\tau$, and thus both $\Sigma^{f}_{{\ell},{h}}(x)=\Sigma^{f}_{{\ell},{h}}(y)=\emptyset$ for all possible choices of $h$ in Algorithm 3. On the other hand, if $m_{\ell}(y)<|y|-4\tau$, then $m_{\ell}(x)\leq|x|-4\tau$. In this case, $\Sigma^{f}_{{\ell},{h}}(x)\neq\emptyset$ if and only if $h=|x|-4\tau$. Since this occurs with probability at most $\delta$, we have $\Pr{\left[{{\mathcal{T}}_{\ell}(x)=1}\right]}\leq\delta$ which completes the proof. ∎
To complete the proof of the lemma, consider the following two cases.
Case 1.
$(x,y)\in G_{\ell}$. In this case, the sets $\Sigma^{f}_{{\ell},{|x|-4\tau}}(x)$ and $\Sigma^{f}_{{\ell},{|y|-4\tau}}(y)$ are nonempty. Observe that for all $u\in\Sigma^{f}_{{\ell},{|x|-4\tau}}(x)$ and $v\in\Sigma^{f}_{{\ell},{|y|-4\tau}}(y)$, we have $|u\cap v|\geq\ell$. Moreover, since $|u|\geq|x|-4\tau$, and $|v|\geq|y|-4\tau$, we must have $|u\setminus(u\cap v)|\leq 4\tau+1$ and $|v\setminus(u\cap v)|\leq 4\tau+1$. Thus, since $u$ and $v$ are $\ell$-stable, we have $|f(u)-f(v)|\leq 8\tau+2\leq 10\tau$. The Laplace mechanism now guarantees that ${\mathcal{A}}_{\ell}(x)\approx_{\varepsilon_{0},0}{\mathcal{A}}_{\ell}(y)$. Since ${\mathcal{T}}_{\ell}(x)\approx_{\varepsilon_{0},\delta_{0}}{\mathcal{T}}_{\ell}(y)$ by B.3, DP composition and postprocessing (2.3 and 2.4) imply that ${\mathcal{P}}_{\ell}(x)\approx_{3\varepsilon_{0},\delta}{\mathcal{P}}_{\ell}(y)$.
Case 2.
$(x,y)\not\in G_{\ell}$. By B.4, both ${\mathcal{T}}_{\ell}(x)\approx_{0,\delta_{0}}\perp\approx_{0,\delta_{0}}{\mathcal{T}}_{\ell}(y)$. Thus, by postprocessing, ${\mathcal{P}}_{\ell}(x)\approx_{0,2\delta_{0}}{\mathcal{P}}_{\ell}(y)$.
Thus, in both cases we have ${\mathcal{P}}_{\ell}(x)\approx_{3\varepsilon_{0},2\delta_{0}}{\mathcal{P}}_{\ell}(y)$. ∎
To see why $\mathcal{W}$ is private, we can simply apply DP composition to the mechanisms $\mathcal{L}$ and ${\mathcal{P}}_{\ell}$. By 2.1, mechanism $\mathcal{L}$ is $(\varepsilon_{0},\delta_{0})$-DP, and thus, by Lemma B.2 and composition the mechanism $\mathcal{W}$ is $(4\varepsilon_{0},3\delta_{0})$-DP.
@@
To see why the accuracy guarantee holds, observe that if $x$ is $\ell$-stable with respect to $f$, then $x=\arg\max\{|v|:v\in\Sigma^{f}_{{\ell},{h-4\tau}}(x)\}$, and hence $\mathcal{W}^{f}(x)$ outputs $f(x)+\text{Lap}{\left({\frac{10\tau}{\varepsilon_{0}}}\right)}$.